OpenStudy (anonymous):

Y varies directly as x. If y=-4 when x=1/2. find y when x=2/3.

OpenStudy (anonymous):

The formula for direct relationships is y=kx. so you have y as -4 and x as 0.5 substitute in the equation and solve to get k as -8. k is your constant in direct relationships. so in your new equation, you now know that k is -8. so the new equation is y=-8x. substitute 2/3 for x and get y=-16/3

OpenStudy (anonymous):

addd the y and x that they gave u?

OpenStudy (anonymous):

first equation second equation y=-4 and x=1/2 y=? x=2/3 -4+1/2=-3.5 y+2/3=-3.5 y=-25/6